OpenStudy (anonymous):

How can I complete the square with this problem: x^2/2 = 4 + x/2

OpenStudy (wolf1728):

For completing the square http://www.1728.org/quadr2.htm x^2/2 = 4 + x/2 On the right, only have the non-x term x^2/2 - x/2 = 4 Divide the equation by the coefficient of x^2 (which is 1/2) x^2 -x = 8 Get coefficient of x and divide it by 2 square it then add it to both sides of the equation -1 divided by 2 = -.5 squared= -.25 x^2 -x -.25 = 7.5 Take the square root of both sides of the equation: (x -.5) = 2.7386127875 x = 3.2386127875

OpenStudy (anonymous):

Thank you.

OpenStudy (wolf1728):

-.25 squared is actually .5 So, x^2 -x +.5 = 8.5 (x-.5)*(x-.5) = 8.5 Okay, there's a little more to do there